### v2

@@ 1 Introduction
The rapid progress in generative models [7, 13] has led the development of virtual try-on methods. Most of the primary approaches for image-based VTON are based on Generative Adversarial Networks [7, 21] (GANs), where they employ a separate warping module to deform the garment to the human body and use GAN generators to blend with the target person [3, 26, 27, 6, 51, 5]. However, such approaches struggle to generate high-quality images and often fail to generalize to different human images, introducing undesirable distortions to the garment. On the other hand, recently, diffusion models [13, 43] demonstrate superior performance in generating genuine images [33, 28, 36, 37] compared to GANs, showing the potential to overcome such challenges. Recent diffusion-based VTON methods [29, 8, 22] utilize the rich generative prior of pretrained text-to-image (T2I) diffusion models [40, 32, 37] to enhance the naturalness of try-on images. To identify the details of garments, they encode the semantics of garments in pseudo-words [29] or use explicit warping networks [8]. However, those methods fall short in preserving the fine-grained details of garments, e.g., patterns, textures, shapes, or colors, which hinders its application to real-world scenarios.
@@ 2 Related Works
As diffusion models have shown great success [13, 44], recent works have studied the application of diffusion models for virtual try-on. TryOnDiffusion [58] proposed a novel architecture that uses two parallel UNets, and showed the capability of diffusion-based virtual try-on by training on the large-scale dataset. Subsequent works considered virtual try-on as an exemplar-based image inpainting problem [53]. They fine-tuned the inpainting diffusion models on virtual try-on dataset [22, 29, 8] to generate high-quality virtual try-on images. However, those methods still suffer from preserving meticulous details of garments, dampening their applications to real-world scenarios.
@@ 3.1 Backgrounds on Diffusion Models
Diffusion models [13, 41] are generative models that consist of a forward process, which gradually adds Gaussian noise to the data, and a reverse process that gradually denoises a random noise to generate a sample. Suppose $\mathbf{x}_{0}$ be a data point (e.g., an image or a latent from the output of autoencoder [37]).Given noise schedules $\{\alpha_{t}\}_{t=1}^{T},\{\sigma_{t}\}_{t=1}^{T}$ for $t=1,\ldots,T$, the forward process at timestep $t$ is given by $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$ is a Gaussian noise. For sufficiently large $\sigma_{T}$, $\mathbf{x}_{T}\sim\mathcal{N}(\alpha_{T}\mathbf{x}_{0},\sigma_{T}^{2}\mathbf{I})$ is indiscernible from pure random Gaussian noise. Then the reverse generative process initializes from $\mathbf{x}_{T}\sim\mathcal{N}(\boldsymbol{0},\sigma_{T}^{2}\mathbf{I})$, and sequentially denoises into $\mathbf{x}_{t}$ so that $\mathbf{x}_{0}$ is distributed according to the data distribution.
@@ 3.2 Proposed Method
While our model is able to capture the fine-details of garments, it often struggles when the person image $\mathbf{x}_{p}$ or the garment image $\mathbf{x}_{g}$ is different from training distribution (e.g., see Fig. 3). To this end, inspired by text-to-image personalization methods [39, 45, 2], we propose an effective customization of IDM–VTON by fine-tuning TryonNet with a pair of garment and person images. It is straightforward to fine-tune IDM–VTON when we have a pair of images $\mathbf{x}_{p}$ and $\mathbf{x}_{g}$, where the person in $\mathbf{x}_{p}$ is wearing the garment in $\mathbf{x}_{g}$. On the other hand, if we only have $\mathbf{x}_{p}$, we segment the garment and white-out backgrounds to obtain $\mathbf{x}_{g}$. Remark that we fine-tune only the attention layers of the up-blocks (i.e., decoder layers) of TryonNet, which empirically works well (e.g., see Table 4).
@@ 4 Experiment
Figure 5: Qualitative comparisons on In-the-Wild dataset. We show generated virtual try-on images on In-the-Wild dataset using IDM–VTON (ours) compared with other methods. IDM–VTON outperforms other methods in generating authentic images and preserving fine-grained details of garment. In particular, customizing IDM–VTON, (i.e., ${}^{\ddagger}$), significantly enhances the image quality and garment fidelity. When applying customization on StableVITON, (i.e., ${}^{\ddagger}$), the improvements are marginal compared to ours. Best viewed in zoomed, color monitor.
@@ 4.1 Experimental Setup
We use SDXL masked inpainting model [46] for our TryonNet, pretrained IP-Adapter [54] for image adapter and UNet of SDXL [33] for our GarmentNet. We train our model using VITON-HD [3] train dataset, which contains $11,647$ person-garment image pairs at $1024\times 768$ resolution. We train our model using Adam [23] optimizer with batch size of $24$ and learning rate of 1e-5 for $130$ epochs. For customization, we fine-tune our model using Adam optimizer with learning rate 1e-6 for $100$ steps.
@@
Table 2: Quantitative results on In-the-Wild dataset. We compare IDM–VTON (ours) with other methods on In-the-Wild dataset to assess the generalization capabilities. We report LPIPS, SSIM and CLIP [34] image similarity scores. For IDM–VTON and StableVITON [22], we further customize models using pair of person-garment images (denoted by ${}^{\ddagger}$). We see that IDM–VTON outperforms other methods, and customized IDM–VTON (i.e., ${}^{\ddagger}$) performs the best.
@@
${}^{\ddagger}$ [22] 0.259 0.733 0.858
${}^{\ddagger}$ (ours) 0.150 0.797 0.909
@@ 4.3 Results on In-the-Wild Dataset
Tab. 2 reports the quantitative comparisons between our IDM–VTON and other methods. We also provide comparisons between IDM–VTON and StableVITON using customization (i.e., ${}^{\ddagger}$ and ${}^{\ddagger}$). We see that IDM–VTON outperforms other methods on all metrics, even outperforming StableVITON with customization. Furthermore, we observe that IDM–VTON with customization significantly outperforms others, showing its generalization capability.
@@ 4.4 Ablative Studies
We conduct an ablation study on the effect of GarmentNet in preserving the fine-grained details of a garment. For comparison, we train another model without using GarmentNet (i.e., using only IP-Adapter to encode garment image). The qualitative results are in Fig. 6. While training with only IP-Adapter is able to generate authentic virtual try-on images that are semantically consistent with the garment images, we see that using GarmentNet significantly improves preserving the identity of the garment (e.g., the graphics of t-shirts). In addition, we conduct a quantitative evaluation on VITON-HD test dataset by measuring LPIPS, SSIM and CLIP image similarity score. In Tab. 4, we see that using GarmentNet quantitatively improves reconstruction scores, as well as image similarity scores, which is aligned with our qualitative results.
@@
During customization, we fine-tune the up-blocks of TryonNet (i.e., decoder layers). Here, we provide an ablation study on the choice of fine-tuning layer for customization. We compare default IDM–VTON (i.e., no customization), fine-tuning all UNet layers, using low rank adaptation (LoRA) [16], and fine-tuning decoder layers. In Tab. 4, we compare the LPIPS, SSIM, and CLIP image similarity scores on In-the-Wild dataset. We see that all customization methods improve reconstruction score compared to default IDM–VTON, while showing comparable performance in CLIP image similarity score. We see that fine-tuning decoder layers perform the best among customization methods, while fine-tuning all UNet layers shows comparable performance. While fine-tuning with LoRA is more efficient, we find it underperforms than fine-tuning decoder layers.
@@ 5 Conclusion
This paper introduces a method that enhances the performance of virtual try-on using generative diffusion models. The virtual try-on technology comes with benefits and pitfalls - the tool could be helpful for users to effectively visualize their look with given garment. However, it requires the users to be responsible for protecting the ownership, and avoid malicious use.
@@
Similar to other inpainting works, IDM–VTON struggles to preserve human attributes on masked regions such as tattoos or skin moles. One can design a method for conditioning those human attributes of human images when generating try-on images. Also, providing comprehensive captions enables fully utilizing the prior knowledge of T2I diffusion models. Exploring broader applications, such as controlling garment generation through textual prompts, is left as future work.
@@ 0.A.3 Detailed captioning of garments
Figure 10: Comparison between IDM–VTON and OOTDiffusion [52] on VITON-HD and DressCode dataset. Both methods are trained on VITON-HD training data. Best viewed in zoomed, color monitor.
@@ Appendix 0.B Comparison with Concurrent Work
Figure 11: Comparison between IDM–VTON and OOTDiffusion [52] on In-the-Wild dataset. We provide visual comparison results on In-the-Wild dataset. Best viewed in zoomed, color monitor.
Table 5: Quantitative results on VITON-HD and DressCode dataset. We additionally compare the metric scores of IDM–VTON (ours) with the concurrent work OOTDiffusion [52].
Dataset VITON-HD DressCode
@@
We additionally compare IDM–VTON with OOTDiffusion [52], which is a concurrent work on virtual try-on task. Both methods utilize an additional UNet to extract features from garment images. However, remark that there are differences between OOTDiffusion and IDM–VTON. Specifically, while OOTDiffusion trains the entire set of parameters of the additional UNet for garment feature extraction (i.e., corresponds to GarmentNet of IDM–VTON), IDM–VTON does not update GarmentNet. We find that pretrained UNet suffices in extracting low-level features, thus our approach is more efficient in terms of cost and memory.
For comparison with OOTDiffusion, we generate try-on images of OOTDiffusion at a resolution of $1024\times 768$, utilizing the model checkpoint trained on VITON-HD dataset, which is available via the official repository [^4] https://github.com/levihsu/OOTDiffusion. Fig. 10 and Fig. 11 present qualitative comparisons of OOTDiffusion and ours on VITON-HD, DressCode and In-the-Wild dataset. As shown in Fig. 10, We see that IDM–VTON outperforms OOTDiffusion in capturing both high-level semantics and low-level details, and generating more authentic images. In particular, we observe notable improvements of IDM–VTON compared to OOTDiffusion on In-the-Wild dataset, which demonstrates the generalization ability of IDM–VTON.
Tab. 5 and Tab. 6 show quantitative comparisons between IDM–VTON and OOTDiffusion on VITON-HD, DressCode, and In-the-Wild datasets. One can notify that IDM–VTON outperforms OOTDiffusion on all metrics including image fidelity (FID), and reconstruction of garments (LPIPS, SSIM and CLIP-I), which verifies our claim.

### v3

@@ 1 Introduction
The rapid progress in generative models [8, 14] has led the development of virtual try-on methods. Most of the primary approaches for image-based VTON are based on Generative Adversarial Networks [8, 22] (GANs), where they employ a separate warping module to deform the garment to the human body and use GAN generators to blend with the target person [3, 27, 29, 7, 54, 6]. However, such approaches struggle to generate high-quality images and often fail to generalize to different human images, introducing undesirable distortions to the garment. On the other hand, recently, diffusion models [14, 46] demonstrate superior performance in generating genuine images [36, 30, 39, 40] compared to GANs, showing the potential to overcome such challenges. Recent diffusion-based VTON methods [31, 9, 23, 4, 28, 35] utilize the rich generative prior of pretrained text-to-image (T2I) diffusion models [43, 34, 40] to enhance the naturalness of try-on images. To identify the details of garments, they encode the semantics of garments in pseudo-words [31] or use explicit warping networks [9]. However, those methods fall short in preserving the fine-grained details of garments, e.g., patterns, textures, shapes, or colors, which hinders its application to real-world scenarios.
@@ 2 Related Works
As diffusion models have shown great success [14, 47], recent works have studied the application of diffusion models for virtual try-on. TryOnDiffusion [61] proposed an architecture that uses two parallel UNets, and showed the capability of diffusion-based virtual try-on by training on the large-scale dataset. Subsequent works considered virtual try-on as an exemplar-based image inpainting problem [56]. They fine-tuned the inpainting diffusion models on virtual try-on dataset [23, 31, 9] to generate high-quality virtual try-on images. However, those methods still suffer from preserving meticulous details of garments, dampening their applications to real-world scenarios.
@@ 3.1 Backgrounds on Diffusion Models
Diffusion models [14, 44] are generative models that consist of a forward process, which gradually adds Gaussian noise to the data, and a reverse process that gradually denoises a random noise to generate a sample. Suppose $\mathbf{x}_{0}$ be a data point (e.g., an image or a latent from the output of autoencoder [40]). Given noise schedules $\{\alpha_{t}\}_{t=1}^{T},\{\sigma_{t}\}_{t=1}^{T}$ for $t=1,\ldots,T$, the forward process at timestep $t$ is given by $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$ is a Gaussian noise. For sufficiently large $\sigma_{T}$, $\mathbf{x}_{T}\sim\mathcal{N}(\alpha_{T}\mathbf{x}_{0},\sigma_{T}^{2}\mathbf{I})$ is indiscernible from pure random Gaussian noise. Then the reverse generative process initializes from $\mathbf{x}_{T}\sim\mathcal{N}(\boldsymbol{0},\sigma_{T}^{2}\mathbf{I})$, and sequentially denoises $\mathbf{x}_{t}$ into $\mathbf{x}_{t-1}$ for each $t$, so that $\mathbf{x}_{0}$ is distributed according to the data distribution.
@@ 3.2 Proposed Method
While our model is able to capture the fine-details of garments, it often struggles when the person image $\mathbf{x}_{p}$ or the garment image $\mathbf{x}_{g}$ is different from training distribution (e.g., see Fig. 3). To this end, inspired by text-to-image personalization methods [42, 48, 2], we propose an effective customization of IDM–VTON by fine-tuning TryonNet with a pair of garment and person images. It is straightforward to fine-tune IDM–VTON when we have a pair of images $\mathbf{x}_{p}$ and $\mathbf{x}_{g}$, where the person in $\mathbf{x}_{p}$ is wearing the garment in $\mathbf{x}_{g}$. On the other hand, if we only have $\mathbf{x}_{p}$, we segment the garment and white-out backgrounds to obtain $\mathbf{x}_{g}$. Remark that we fine-tune only the attention layers of the up-blocks (i.e., decoder layers) of TryonNet, which empirically works well (e.g., see Table LABEL:tab:adapt).
@@ 4 Experiment
Figure 5: Qualitative comparisons on In-the-Wild dataset. We show generated virtual try-on images on In-the-Wild dataset using IDM–VTON (ours) compared with other methods. IDM–VTON outperforms other methods in generating authentic images and preserving fine-grained details of garment. In particular, customizing IDM–VTON, (i.e., IDM–VTON), significantly enhances the image quality and garment fidelity. When applying customization on StableVITON, (i.e., StableVITON), the improvements are marginal compared to ours. Best viewed in zoomed, color monitor.
@@ 4.1 Experimental Setup
We use SDXL inpainting model [49] for our TryonNet, pretrained IP-Adapter [57] for image adapter and UNet of SDXL [36] for our GarmentNet. We train our model using VITON-HD [3] train dataset, which contains $11,647$ person-garment image pairs at $1024\times 768$ resolution. We train our model using Adam [24] optimizer with batch size of $24$ and learning rate of 1e-5 for $130$ epochs. For customization, we fine-tune our model using Adam optimizer with learning rate 1e-6 for $100$ steps.
@@
Table 2: Quantitative results on In-the-Wild dataset. We compare IDM–VTON (ours) with other methods on In-the-Wild dataset to assess the generalization capabilities. We report LPIPS, SSIM and CLIP [37] image similarity scores. For IDM–VTON and StableVITON [23], we further customize models using a single pair of person-garment images (denoted by ). We see that IDM–VTON outperforms other methods, and customized IDM–VTON (i.e., IDM–VTON) performs the best.
@@
Stable-VITON [23] 0.259 0.733 0.858
IDM–VTON (ours) 0.150 0.797 0.909
@@ 4.3 Results on In-the-Wild Dataset
Tab. 2 reports the quantitative comparisons between our IDM–VTON and other methods. We also provide comparisons between IDM–VTON and StableVITON using customization (i.e., IDM–VTON and StableVITON). We see that IDM–VTON outperforms other methods on all metrics, even outperforming StableVITON with customization. Furthermore, we observe that IDM–VTON with customization significantly outperforms others, showing its generalization capability.
@@ 4.4 Ablative Studies
We conduct an ablation study on the effect of GarmentNet in preserving the fine-grained details of a garment. For comparison, we train another model without using GarmentNet (i.e., using only IP-Adapter to encode garment image). The qualitative results are in Fig. 6. While training with only IP-Adapter is able to generate authentic virtual try-on images that are semantically consistent with the garment images, we see that using GarmentNet significantly improves preserving the identity of the garment (e.g., the graphics of t-shirts). In addition, we conduct a quantitative evaluation on VITON-HD test dataset by measuring LPIPS, SSIM and CLIP image similarity score. In Tab. LABEL:tab:ablunet, we see that using GarmentNet quantitatively improves reconstruction scores, as well as image similarity scores, which is aligned with our qualitative results.
@@
During customization, we fine-tune the up-blocks of TryonNet (i.e., decoder layers). Here, we provide an ablation study on the choice of fine-tuning layer for customization. We compare default IDM–VTON (i.e., no customization), fine-tuning all UNet layers, using low rank adaptation (LoRA) [17], and fine-tuning decoder layers. In Tab. LABEL:tab:adapt, we see that all customization methods improve reconstruction score (LPIPS, SSIM) compared to default IDM–VTON, while showing comparable performance in CLIP image similarity score. We see that fine-tuning decoder layers perform the best among customization methods, while fine-tuning all UNet layers shows comparable performance. While fine-tuning with LoRA is more efficient, we find it underperforms than fine-tuning decoder layers.
@@ 5 Conclusion
The virtual try-on technology comes with benefits and pitfalls - the tool could be helpful for users to effectively visualize their look with given garment. However, it requires the users to be responsible for protecting the ownership, and avoid malicious use.
@@
IDM–VTON struggles to preserve human attributes on masked regions such as tattoos or skin moles. One can design a method for conditioning those human attributes of when generating try-on images. Also, while we exploit the prior knowledge of T2I diffusion models by providing comprehensive captions, exploring broader applications to fully utilize the prior, such as controlling garment generation through textual prompts, is left as future work.
Acknowledgement
This work was supported by Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No.2019-0-00075 Artificial Intelligence Graduate School Program(KAIST); No.RS-2021-II212068, Artificial Intelligence Innovation Hub).
@@ 0.A.3 Detailed captioning of garments
Figure 10: Comparison between IDM–VTON and OOTDiffusion [55] on VITON-HD and DressCode dataset. Both methods are trained on VITON-HD training dataset. Best viewed in zoomed, color monitor.
@@ Appendix 0.B Comparison with Concurrent Work
Figure 11: Comparison between IDM–VTON and OOTDiffusion [55] on In-the-Wild dataset. We provide visual comparison results on In-the-Wild dataset. Both methods are trained on VITON-HD training dataset. Best viewed in zoomed, color monitor.
Table 5: Quantitative results of models trained on VITON-HD training dataset and evaluated on VITON-HD and DressCode (upper body) test dataset. We additionally compare the metric scores of IDM–VTON (ours) with the concurrent work OOTDiffusion [55].
Train / Test VITON-HD / VITON-HD VITON-HD / DressCode
@@
Table 6: Quantitative results of models trained on DressCode training dataset and evaluated on VITON-HD and DressCode test dataset. We additionally compare the metric scores of IDM–VTON (ours) with the concurrent work OOTDiffusion [55].
Train / Test DressCode / DressCode DressCode / VITON-HD
OOTDiffusion [55] 0.088 0.885 4.18 0.851 0.205 0.788 57.87 0.677
IDM–VTON (ours) 0.072 0.902 3.62 0.913 0.129 0.859 12.57 0.839
@@
We additionally compare IDM–VTON with OOTDiffusion [55], which is a concurrent work on virtual try-on task. We compare each model trained on VITON-HD and DressCode datasets. We use publicly available model checkpoints for generating try-on images of OOTDiffusion [^4] https://github.com/levihsu/OOTDiffusion.
Fig. 10 and Fig. 11 present qualitative comparisons of OOTDiffusion and ours on VITON-HD, DressCode and In-the-Wild dataset. As shown in Fig. 10, we see that IDM–VTON outperforms OOTDiffusion in capturing both high-level semantics and low-level details, and generating more authentic images. In particular, we observe notable improvements of IDM–VTON compared to OOTDiffusion on In-the-Wild dataset, which demonstrates the generalization ability of IDM–VTON.
Tab. 5, Tab. 6 and Tab. 7 show quantitative comparisons between IDM–VTON and OOTDiffusion on VITON-HD, DressCode, and In-the-Wild datasets. One can notify that IDM–VTON outperforms OOTDiffusion on all metrics including image fidelity (FID), and reconstruction of garments (LPIPS, SSIM and CLIP-I), which verifies our claim.